You wish to make a 0.398 M hydroiodic acid solution from a stock solution of 3.00 M hydroiodic acid. How much concentrated acid must you add to obtain a total volume of 100 mL of the dilute solution?
1
Expert's answer
2019-03-18T08:12:57-0400

"C_m= \frac{\nu}{V}"

Cm - molarity or molar concentration of the solution

"\nu" - amount of the solute

V - volume of the solution


"\nu(HI)=const=C_{m1} \cdot V_1=C_{m2} \cdot V_2"

Cm1 - initial molarity of the solution

V1 - initial volume of the solution

Cm2 - final molarity of the solution

V2 - final volume of the solution

"V_1=\frac{C_{m2} \cdot V_2}{C_{m1}}"

"V_1=\frac{0.398 \; M \cdot 100.0 \; mL}{3.00 \; M} \approx 13.27 \; mL"


Answer: 13.27 mL.